OpenStudy (anonymous):

Which of the following is equal to (6-m)? A. m-6 B. -(m-6) C. -(m+6)

OpenStudy (anonymous):

I'd say A would be equal to that.

OpenStudy (anonymous):

What happens when you distribute the negative sign, in (B)?

OpenStudy (anonymous):

m-6 is not the same as 6-m. That would be like saying 1-100 is the same as 100-1.

OpenStudy (anonymous):

-(m-6) = -m --6 = -m+6 = 6-m B